OpenStudy (anonymous):

If sin α = − 5/13 and tan α > 0, find the exact value of sin(α − 5π/3).

OpenStudy (anonymous):

you gotta use the subtraction angle formula for sine do you know it?

OpenStudy (anonymous):

if the answer is "no i have no clue" that is fine, i will write it for you

OpenStudy (anonymous):

btw this problem is not hard, but it is a pain in the ...

OpenStudy (anonymous):

is it sin(a - b) = (sina)(cosb) -(cosa)(sinb) ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so what is missing is the \(\cos(\alpha)\)

OpenStudy (anonymous):

all the rest you know

OpenStudy (anonymous):

if the sin of 5pi/3 = sqrt-(3)/2, is the cos(a) = 1/2 ?

OpenStudy (anonymous):

No I don't! Help me set the problem up??

OpenStudy (anonymous):

\[\sin(\frac{5\pi}{3})=-\frac{\sqrt3}{2}\] \[\cos(\frac{5\pi}{3})=\frac{1}{2}\] \[\sin(\alpha)=-\frac{5}{13}\] you have 3 of the four numbers

OpenStudy (anonymous):

the one that is missing is \[\cos(\alpha)\] which you get by drawing a triangle

OpenStudy (anonymous):



OpenStudy (anonymous):

all that you need is the missing side do you know what it is?

OpenStudy (calculusfunctions):

First you're told that sine is negative and tangent is positive. This happens in the third quadrant.  Now using Pythagorean theorem, find the third side length of this triangle.

OpenStudy (anonymous):

The third side = 12?

OpenStudy (calculusfunctions):

Then use the idntity\[\sin (A -B)=\sin A \cos B -\cos A \sin B\]

OpenStudy (calculusfunctions):

Yes but it's -12 (the third side because it extends on the negative x-axis). Understand @lizbens8 ??

OpenStudy (anonymous):

you can think of it as 12 or -12, but \[\cos(\alpha)=-\frac{12}{13}\]

OpenStudy (anonymous):

now plug in the numbers is all that is left to do you have all four numbers that you need

OpenStudy (anonymous):

\[\sin(\alpha-\beta) = (5\pi/3)(12/13) - (1/2 - \sin \beta) ??\]

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

just put the numbers in the formula

OpenStudy (anonymous):

\[\sin (\alpha -\beta)=\sin (\alpha)\cos (\beta) -\cos (\alpha) \sin (\beta)\] you know all four numbers

OpenStudy (calculusfunctions):

NO. You mean\[\sin (\alpha - \frac{ 5\pi }{ 3})=\sin \alpha \cos \frac{ 5\pi }{ 3 }-\cos \alpha \sin \frac{ 5\pi }{ 3 }\]

OpenStudy (anonymous):

\[\sin(\frac{5\pi}{3})=-\frac{\sqrt3}{2}\\ \sin(\alpha)=-\frac{5}{13}\\ \cos(\frac{5\pi}{3})=\frac{1}{2}\\ \cos(\alpha)=-\frac{12}{13}\] as you said earlier

OpenStudy (anonymous):

put the numbers in the right spot is all

OpenStudy (calculusfunctions):

What doesn't make sense is that the angle 5pi/3 is in the fourth quadrant and in the question you wrote that tangent of alpha > 0. I think either you copied it incorrectly or the textbook had it wrong but either way it should be tangent < 0.

OpenStudy (anonymous):

Okay so \[-5 - 12\sqrt{3}/26\] is what I got

OpenStudy (anonymous):

i didn't do it but it looks reasonable want me to check?

OpenStudy (anonymous):

\[-\frac{5}{13}\times \frac{1}{2}-(-\frac{12}{13})\times (-\frac{\sqrt3}{2})\] is the first step

OpenStudy (anonymous):

so yeah, what you got

OpenStudy (anonymous):

I checked on my homework answers and that was it! Thank you!

OpenStudy (anonymous):

yw, hope it was more or less clear